OpenStudy (marco26):

10. The sides of the triangle are 40 cm, 50 cm, and 60 cm. How far is the point of intersection of the perpendicular bisector of the sides of the triangle to any of its vertex?

OpenStudy (anonymous):

it is radius of circumcircle. \[R=\frac{ abc }{ 4~ area~ of~ \triangle }\]

OpenStudy (anonymous):

area of triangle\[=\sqrt{s \left( s-a \right)\left( s-b \right)\left( s-c \right)}\] where a,b,c are sides of triangle. and 2s=a+b+c

OpenStudy (marco26):

ok thanks, I get it now :)

OpenStudy (anonymous):

yw